OpenStudy (anonymous):

The perimeter of a rectangle is 28ft. The length is 2ft. longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. Write the first equation 2x+2y=? Write the second equation x=y+? What is the length? What is the width?

hartnn (hartnn):

the perimeter = 28 so 2x+2y = 28 length is 2 feet longer, so if we add 2 to width,it equals length so x=y+2 got it?

OpenStudy (anonymous):

so now i just solve the equation? to get the length and width?

hartnn (hartnn):

yup, can u ? or need the first step ?

OpenStudy (anonymous):

i think so...let me check..

OpenStudy (anonymous):

so 2x+2y=28 which is 14? so is that the length

hartnn (hartnn):

x+y = 14 and x-y = 2 now u can add these 2 . tell me what u get then.....

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i got the answer 14 and 2...y do i need to add is those two no the answer of length and width

hartnn (hartnn):

i will add for u x+y+x-y = 14+2 ok? so 2x+0 = 16 so x= 16/2 = ??

hartnn (hartnn):

14 and 2 are not length and width, they are x and y , u need to solve those equations to get x and y

hartnn (hartnn):

after solving as above u get x=16/2 = 8 so y= x-2= 6 now 8 and 6 are your length and width respectively, got it?

OpenStudy (anonymous):

yes i think so..you were very helpful thank you so much!

hartnn (hartnn):

welcome :) make sure u understand it .....